OpenStudy (anonymous):

3 square root of 64.... the three is a little three in the check area of the square root sign

OpenStudy (anonymous):

The little 3 means that you're trying to find the cube root of 64:\[\sqrt[3]{64}\] In other words, you want a number x such that \[ x^3 = 64 \]

OpenStudy (anonymous):

thats what it looks like!

OpenStudy (anonymous):

Try testing out numbers... \[ 2^3 = 2 * 2 * 2 = 8 < 64\] \[ 5^3 = 5 * 5 * 5 = 125 >64 \] So the x you're looking for is between 2 and 5...

OpenStudy (anonymous):

Why don't you cube 3 and 4 to see which ones correct?

OpenStudy (anonymous):

\[^3\sqrt{64}\] it says to find the value of that

OpenStudy (anonymous):

Right, but \[\sqrt[3]{64}=x\] can also be written as \[64 = x^3\] (because roots and exponents are opposites) Does that make sense?

OpenStudy (anonymous):

The cubed root of 64 = 4. 4x4x4=64